OpenStudy (anonymous):

no matter what i do i keep getting decimals someone please help me.One-fifth the sum of three consecutive even integers is −6. Find the largest integer.

OpenStudy (turingtest):

how did you set up the problem?

OpenStudy (anonymous):

1/5(3x-6) 0.6x-0.2=-6 +0.2 =0.6x/0.6= -5.8/0.6

OpenStudy (anonymous):

where did i go wrong?

OpenStudy (radar):